# An insight on some properties of high order nonstandard linear multistep methods

### v1

@@ Abstract
In this paper, nonstandard multistep methods are observed. It is shown that under some (sufficient and necessary) conditions, these methods attain the same order as their standard counterparts - to prove this statement, a nonstandard version of Taylor’s series is constructed. The preservation of some qualitative properties (boundedness, a weak form of monotonicity, and the linear combination of the components) is also proven for all step sizes. The methods are applied to a one-dimensional equation and a system of equations, in which the numerical experiments confirm the theoretical results.
@@ 1 Introduction
Two key aspects of the numerical modeling of differential equations have always been the high accuracy of the methods along with the preservation of important features of the initial continuous problem. It has been known that the application of high-order methods does not necessarily imply the latter (see e.g. [1], along with the specific cases of multistep methods: [2, 3]).
During the previous decades, numerous different approaches have been taken to construct methods with such properties. One such approach, which has had an increasing prevalence in recent years, has been the approach of nonstandard methods: only last year, multiple articles applied these ideas to epidemic problems [4, 5, 6, 7, 8], to equations of population dynamics [9, 10, 11, 12], to models in microbiology and chemistry [13, 14], to problems in finance [15, 16] and also to equations with derivatives of fractional order [17, 18, 19, 20], just to name a few (for a comprehensive list of earlier applications, the author recommends the remarkably insightful survey articles of Patidar [21, 22] and the references therein). The main advantage of the use of such methods is that they are customarily constructed in a way that some desired properties of the original, continuous model are preserved for any choice of step size.
@@
Earliest examples include the application of these ideas to the Yee algorithm [23, 24] and the observation of the second-order nonstandard Runge-Kutta method [25, 26]. Nevertheless, nowadays one leading school of thought is the use of theta methods, which can achieve an order of two [27, 28], can be combined with second order Runge-Kutta methods [29], and can also be applied to reaction-diffusion [30, 31] and singularly perturbed problems [32].
Another possibility is the application of Richardson extrapolation [33], which can also be applied to population models [34]. There are also approaches involving the use of a more complicated form of the denominator function [35, 36, 37, 38], while other ideas incorporate innovative modifications of the right-hand side of the method [39, 40]. Other works include the general order conditions for one-step nonstandard methods applied to some special types of equations [41, 42] or methods with variable step sizes [43]. These ideas have also been applied to several real-life applications [13, 44, 45, 10].
It should also be mentioned that the application of strongly stable methods in the context of nonstandard methods (which will be the core idea in the preservation of the properties in our case) has appeared in some previous works: some have combined the nonstandard methods with total variation diminishing (TVD) schemes [46] and Runge-Kutta methods written in a strongly stability preserving (SSP) form [47] - the authors in [48] even got the order condition for second order methods akin to the one proven in Section 3.1 (other similar instances include [35] and [38]).
One article to highlight is the work of Dang and Hoang [49] involving nonstandard Runge-Kutta methods. In this, it is shown that the nonstandard versions of these schemes have the same order as their standard counterparts, and (under some conditions) they also preserve the properties of the initial differential equation. It should also be pointed out that the condition of convergence in that work corresponds strongly to the one that is proven in this present article in Section 3.1 (however, their condition is only sufficient while the one in this paper is also necessary).
In this paper, we focus on nonstandard multistep methods. Such schemes were introduced in [50] by Anguelov and Lubuma. However, due to the issues mentioned in Section 2, these ideas were abandoned and not considered by other authors. This work aims to establish that such issues can be resolved by imposing some additional conditions on the denominator function. In Section 2, we state the most important definitions and core concepts. The main results are presented in Section 3 with their proofs detailed in Section 4. Two possible applications are mentioned in Section 5 with several numerical solutions illustrating the effectiveness of the methods. The article concludes with some possible extensions in Section 6.
@@ 2 Preliminaries
In this paper, we observe an initial problem of an autonomous ordinary differential equation given in the form
$\begin{cases}y^{\prime}(t)&=f(y(t)),\\ y(t_{0})&=\tilde{y},\end{cases}$ (1)
where $y:[t_{0},\;T]\rightarrow\mathbb{R}^{m}$ ($[t_{0},\;T]\subset\mathbb{R}$, $m\in\mathbb{Z}^{+}$), $y(t)=\left(y_{1}(t),\;y_{2}(t),\;\dots,\;y_{m}(t)\right)^{T}$, $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ and $\tilde{y}\in\mathbb{R}^{m}$. We assume that (1) has a unique solution (this can be ensured if $f$ has the Lipschitz property).
@@
where $D_{\Delta t}u^{n}$ is the approximation of the derivative $y^{\prime}(t)$ of the exact solution $y(t)$, and $F_{\Delta t}(f;u^{n})$ approximates the right-hand side of the differential equation. Later, the notation $F_{\Delta t}(u^{n})$ is also used.
@@
which is beneficial when observing the preservation of certain qualitative properties of the method. (We can also include implicit methods if the summation starts at $j=0$.)
In [50], Anguelov and Lubuma proposed the nonstandard version of the previous (’standard’) method (in line with the nonstandard modeling rules of Mickens, see e.g. [1]), which takes the form
@@
In that work, the consistency and the zero-stability of the non-standard method were proven. Namely,
@@
If (3) represents a standard finite difference scheme that is consistent and zero-stable, then any corresponding nonstandard finite difference scheme (5) (that is, with the same $\alpha_{j}$ and $\beta_{j}$ constants) is necessarily consistent. Furthermore, scheme (5) is zero-stable provided that the operator $F_{\Delta t}$ satisfies, for some $M>0$ independent of $h$ and for any bounded sequences $(u_{n})$ and $(\tilde{u}_{n})$ in $\mathbb{R}^{n}$, the Lipschitz condition
@@
In [50], Anguelov and Lubuma observed the elementary stability properties of the linear multistep methods. A numerical method is called elementary stable if the numerical scheme has the same equilibrium points as the continuous model, and their stability is also the same. It was shown that under some mild conditions (with the absolute stability region of the numerical method being bounded, and the original differential equation having at least one stable equilibrium), the ’classical’ linear multistep methods are not elementary stable when their order is greater than one. Because of this, they focused on one-step methods.
In the subsequent section, we demonstrate that under some conditions, the nonstandard linear multistep method (5) attains the same order as the ’standard’ counterpart (3), and it also preserves some key qualitative properties: namely, positivity, a weaker form of monotonicity, and the linear combination of the elements of $u^{n}$.
@@ 3 Main results
In this section, we present the main results of the paper. First, the order of the nonstandard multistep methods is observed, and then the preservation of qualitative properties is shown. The proofs of these statements can be found in Section 4.
@@ 3.1 The order of nonstandard multistep methods
Before presenting the theorem about the order of nonstandard linear multistep methods, the lemma used in the theorem’s proof is highlighted. The reason for this is that the condition required in the statement of the lemma has substantial implications to be discussed later.
@@
Note that if $\varphi(x)=x$, then we get the ’usual’ form of Taylor’s theorem. Moreover, one could rephrase Lemma 1 in the following way: formula (8) holds if and only if the first $(n+1)$-many terms of the Taylor expansions (around zero) of the functions $\varphi(x)$ and $f(x)=x$ are the same. In Section 3.3 we show several possible choices of $\varphi$ for which condition (C1) holds for different values of $n$, and these choices are compared in Section 5.
@@
It is worth mentioning here that one can generalize the previous lemma in the following way: let us assume that (other than the smoothness assumption in the statement of Lemma 1) $\varphi(0)=0$, $\varphi(x)>0$ as $x>0$, and $\varphi^{\prime}(x)\neq 0$ as $x>0$. Then, a similar form can be proven, namely
@@
where $\xi\in(a,x)$. (Note that for $\varphi(x)=x$, this is Taylor’s theorem with the Lagrange form of the remainder.) Nevertheless, because of the last term of (9), this form is not beneficial in our case.
@@
Let us assume that for equation (1) with a smooth solution $y\in C^{p+1}([t_{0},T])$, the ’standard’ linear multistep method (3) is convergent of order $p$. Let us assume that for function $\varphi(x)$, it is $(p+1)$-times continuously differentiable, and neither $\varphi$ nor its derivatives vanish for $x>0$. Then, the nonstandard linear multistep method (5) is convergent of order $p$ if and only if condition (C1) of Lemma 1 holds with $n=p$.
@@ 3.2 Preservation of qualitative properties
In this section, we observe the preservation of some qualitative properties of the original differential equation (1). In all three cases, we consider the Forward Euler method applied to equation (1) with timestep $\Delta t_{FE}$, and denote the resulting numerical scheme by $v^{n}=(v_{1}^{n},\;v_{2}^{n},\;\dots,\;v_{m}^{n})^{T}$.
Then, we assume that this method preserves some qualitative property under some conditions for the timestep, and then observe the same property for the explicit nonstandard multistep method with $s$-many steps written in the form
@@
[…]
@@
Another theorem can be stated if we consider the monotone property of the solution. Let us assume that the component $y_{k}(t)$ of the solution of equation (1) is monotone increasing for $t\in[t_{0},T]$ if $\tilde{y}\in(A,B)$ element-wise ($A,B\in\mathbb{R},\;A<B$, where we also allow that $A=-\infty$ or $B=\infty$).
@@
  • For the starting values of the multistep method (10), we have $u_{k}^{0},\;u_{k}^{1},\;\dots,\;u_{k}^{s-1}\in(A,B)$.
@@
Note that property (11) is only a necessary, but not sufficient condition for the monotone increase of numerical solution $u_{k}^{n}$. Numerical experiments indicate that the numerical scheme can indeed preserve the monotone property if the bound for $\varphi(x)$ is even smaller than $\mathcal{C}B_{FE}$, but for $\varphi(x)\leq\mathcal{C}B_{FE}$, only inequality (11) holds (see the tests in Section 5). The determination of a bound under which the monotonicity is also preserved could be a topic of further research.
@@
[…]

### v2

@@ Abstract
In this paper, nonstandard multistep methods are considered. It is shown that under some (sufficient and necessary) conditions, these methods attain the same order as their standard counterparts - to prove this statement, a nonstandard version of Taylor’s series is constructed. The preservation of some qualitative properties (boundedness, the linear combination of the components , and a property similar to monotonicity) is also proven for all step sizes. The methods are applied to a one-dimensional equation and a system of equations, in which the numerical experiments confirm the theoretical results.
@@ 1 Introduction
Two key aspects of the numerical modeling of differential equations have always been the high accuracy of the methods along with the preservation of important features of the initial continuous problem. It has been known that the application of high-order methods does not necessarily imply the latter: see e.g. [1], along with the specific cases of multistep methods [2, 3].
During the previous decades, numerous different approaches have been taken to construct methods with such properties. One of such approaches, which has had an increasing prevalence in recent years, has been the approach of nonstandard methods: only last year, multiple articles applied these ideas to epidemic problems [4, 5, 6, 7, 8], to equations of population dynamics [9, 10, 11, 12], to models in microbiology and chemistry [13, 14], to problems in finance [15, 16] and also to equations with derivatives of fractional order [17, 18, 19, 20], just to name a few. For a comprehensive list of earlier applications, the author recommends the remarkably insightful survey articles of Patidar [21, 22] and the references therein. The main advantage of the use of such methods is that they are customarily constructed in a way that some desired properties of the original, continuous model are preserved for any choice of step size.
@@
Earliest examples include the application of these ideas to the Yee algorithm [23, 24] and the analysis of the second-order nonstandard Runge-Kutta method [25, 26]. Nevertheless, nowadays one leading school of thought is the use of theta methods, which can achieve an order of two [27, 28], can be combined with second order Runge-Kutta methods [29], and can also be applied to reaction-diffusion [30, 31] and singularly perturbed problems [32].
Another possibility is the application of Richardson extrapolation [33], which can also be applied to population models [34]. There are also approaches involving the use of a more complicated form of the denominator function [35, 36, 37, 38, 39, 40], while other ideas incorporate innovative modifications of the right-hand side of the equation [41, 42]. Other works include the general order conditions for one-step nonstandard methods applied to some special types of equations [43, 44] or methods with variable step sizes [45]. These ideas have also been applied to several real-life applications [13, 46, 47, 10].
It should also be mentioned that the core idea in the preservation of the properties in the case of nonstandard multistep methods, namely, the application of strongly stable methods in the context of nonstandard methods, has appeared in some previous works: some have combined the nonstandard methods with total variation diminishing (TVD) schemes [48] and Runge-Kutta methods written in a strongly stability preserving (SSP) form [49] - the authors in [50] even got the order condition for second order methods akin to the one proven in Section 3.1 (other similar instances include [35] and [38]).
One article to highlight is the work of Dang and Hoang [51] involving nonstandard Runge-Kutta methods. In this, it is shown that the nonstandard versions of these schemes have the same order as their standard counterparts, and (under some conditions) they also preserve the properties of the initial differential equation. It should also be pointed out that the condition of convergence in that work corresponds strongly to the one that is proven in this present article in Section 3.1. However, their condition is only sufficient while the one in this paper is also necessary.
In this paper, we focus on nonstandard multistep methods. Such schemes were introduced in [52] by Anguelov and Lubuma. However, due to the issues mentioned in Section 2, these ideas were abandoned and not considered by other authors. This work aims to establish that such issues can be resolved by imposing some additional conditions on the denominator function. In Section 2, we state the most important definitions and core concepts. The main results are presented in Section 3 with their proofs detailed in Section 4. Two possible applications are mentioned in Section 5 with several numerical simulations illustrating the effectiveness of the methods. The article concludes with some possible extensions in Section 6.
@@ 2 Preliminaries
In this paper, we analyze an initial problem of an autonomous ordinary differential equation given in the form
@@
where $y:[t_{0},\;T]\rightarrow\mathbb{R}^{m}$ ($t_{0},\;T]\subset\mathbb{R}$, $m\in\mathbb{Z}^{+}$), $y(t)=\left(y_{1}(t),\;y_{2}(t),\;\dots,\;y_{m}(t)\right)^{T}$, $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ and $\tilde{y}\in\mathbb{R}^{m}$. We assume that (1) has a unique global solution : this can be ensured if $f$ has the global Lipschitz property , but other, weaker conditions can also be assumed.
To get an approximate solution of (1), we split our time-interval $t_{0},\;T]$ with equidistant points $\{t_{n}:=t_{0}+n\Delta t\}_{n=0,\;1,\;\dots,\;\mathcal{N}}$, where $t_{\mathcal{N}}=T$ and $\Delta t>0$ is the step-size. For the approximations of $y(t)$ at point $t=t_{n}$, we use the notation $u^{n}\approx y(t_{n})$. In line with the notations of [52], the resulting numerical scheme can be written in the form
@@
where $D_{\Delta t}u^{n}$ is the approximation of the derivative of the exact solution $y(t)$ at $t_{n}$, and $F_{\Delta t}(f;u^{n})$ approximates the right-hand side of the differential equation. Later, the notation $F_{\Delta t}(u^{n})$ is also used.
@@
which is beneficial when analyzing the preservation of certain qualitative properties of the method. We can also include implicit methods if the summation starts at $j=0$.
In [52], in line with the nonstandard modeling rules of Mickens, see e.g. [1], Anguelov and Lubuma proposed the nonstandard version of the previous ( “standard”) method, which takes the form
@@
It should be noted that although method (5) is implicit, in this paper we only consider the explicit case.
Moreover, another important remark is that from the two usual guiding principles of nonstandard methods, only one of them (the use of a denominator function) is taken into account, while the application of nonlocal approximations in not considered. The case when both of them are implemented could be a topic of further research.
In [52], the consistency and the zero-stability of method (5) were proven. Namely,
@@
If (3) represents a standard finite difference scheme that is consistent and zero-stable, then any corresponding nonstandard finite difference scheme (5) (that is, with the same $\alpha_{j}$ and $\beta_{j}$ constants) is necessarily consistent. Furthermore, scheme (5) is zero-stable provided that the operator $F_{\Delta t}$ (as defined before) satisfies, for some $M>0$ independent of $h$ and for any bounded sequences $(u_{n})$ and $(\tilde{u}_{n})$ in $\mathbb{R}^{n}$, the Lipschitz condition
@@
In [52], Anguelov and Lubuma analyzed the elementary stability properties of the linear multistep methods. A numerical method is called elementary stable if the numerical scheme has the same equilibrium points as the continuous model, and their stability is also the same. It was shown that under some mild conditions (with the absolute stability region of the numerical method being bounded, and the original differential equation having at least one stable equilibrium), the “classical” linear multistep methods are not elementary stable when their order is greater than one. Because of this, they focused on one-step methods.
In the subsequent section, we demonstrate that under some conditions, the nonstandard linear multistep method (5) attains the same order as the “standard” counterpart (3), and it also preserves some key qualitative properties: namely, boundedness, a property similar to monotonicity, and the linear combination of the elements of $u^{n}$.
@@ 3 Main results
In this section, we present the main results of the paper. First, the order of the nonstandard multistep methods is analyzed, and then the preservation of qualitative properties is shown. The proofs of these statements can be found in Section 4.
@@ 3.1 The order of nonstandard multistep methods
Before presenting the theorem about the order of nonstandard linear multistep methods, the lemma used in the theorem’s proof is highlighted. In the proof of the convergence of the nonstandard multistep methods, this nonstandard form of the usual Taylor expansion will be used.
@@
Note that if $\varphi(x)=x$, then we get the “usual” form of Taylor’s theorem. Moreover, one could rephrase Lemma 1 in the following way: formula (8) holds if and only if the first $(n+1)$-many terms of the Taylor expansions (around zero) of the functions $\varphi(x)$ and $f(x)=x$ are the same. In Section 3.3 we show several possible choices of $\varphi$ for which condition (C1) holds for different values of $n$, and these choices are compared in Section 5.
@@
It is worth mentioning here that one can generalize the previous lemma in the following way: let us assume that , other than the smoothness assumption in the statement of Lemma 1 , $\varphi(0)=0$, $\varphi(x)>0$ as $x>0$, and $\varphi^{\prime}(x)\neq 0$ as $x>0$. Then, by using the intermediate value theorem, a similar form can be proven , namely
@@
where $\xi\in(a,x)$. Note that for $\varphi(x)=x$, this is Taylor’s theorem with the Lagrange form of the remainder. Nevertheless, because of the last sum of (9), this form is not beneficial in our case.
@@
Let us assume that for equation (1) with a smooth solution $y\in C^{p+1}([t_{0},T])$, the “standard” linear multistep method (3) is convergent of order $p$. Let us assume that for function $\varphi(x)$, it is $(p+1)$-times continuously differentiable, and neither $\varphi$ nor its derivatives vanish for $x>0$. Then, the nonstandard linear multistep method (5) is convergent of order $p$ if and only if condition (C1) of Lemma 1 holds with $n=p$.
Remark 3.
Note that Theorem 2 can also be proven without the use of Lemma 1, and directly by the fact that condition (C1) is equivalent to
  1. (C3)
    $\varphi(\Delta t)=\Delta t+O((\Delta t)^{n+1})$
(see the proof of Lemma 1). However, the author believes that Lemma 1 gives a further insight into the necessity of condition $\varphi(\Delta t)=\Delta t+O((\Delta t)^{n+1})$: namely, the nonstandard Taylor expansion works if and only if (C1) (and consequently, (C3)) is true, and this nonstandard expansion ensures the convergence of the nonstandard methods (similarly as the “standard” Taylor expansion in the case of “standard” methods). In other words, by using Lemma 1, one can get a sufficient and necessary condition for the order of the method, and the reason for the necessity of these conditions is illustrated by Lemma 1.
Remark 4.
It should be mentioned here that the absolute stability regions (see e.g. [56], Ch. V.1 Def. 1.1.) of nonstandard multistep methods coincide with the same regions of the standard ones - the main reason for this is that the corresponding characteristic equation
$\rho(\zeta)-\mu\sigma(\zeta)=0$
is the same as in the nonstandard case, the only difference being that the variable $\mu$ is now defined as $\mu=\varphi(\Delta t)\lambda$ instead of $\mu=\Delta t\lambda$, but because there are no restrictions on the range of $\lambda\in\mathbb{C}$, variable $\mu$ can also attain any complex value, even if $\varphi$ is bounded (see the next Section). Because of this, the second Dahlquist barrier (see [57], along with [56] and [58]) of standard methods is also true in this, nonstandard setting.
@@ 3.2 Preservation of qualitative properties
In this section, we consider the preservation of some qualitative properties of the original differential equation (1). In all three cases, we consider the Forward Euler method applied to equation (1) with timestep $\Delta t_{FE}$, and denote the resulting numerical solution by $v^{n}=(v_{1}^{n},\;v_{2}^{n},\;\dots,\;v_{m}^{n})^{T}$.
Then, we assume that this method preserves some qualitative property under some conditions for the timestep, and then analyze the same property for the explicit nonstandard multistep method with $s$-many steps written in the form
@@
Remark 5.
It should be noted here that the derivation of the bound $\Delta t_{FE}$ can be challenging sometimes, and a general algorithm for its calculation might be complicated. However, for simple equations it can be determined: for example, if the solution of the equation is monotone. Two calculations can be found among the numerical experiments in Section 5.
Remark 6.
A topic that has been omitted until this point is the choice of the initial values of the numerical method. In [51], it was shown that nonstandard versions of Runge-Kutta methods (i.e. ones, where the usual $\Delta t$ term is interchanged with a $\varphi(\Delta t)$ term), have (at least) the same order $p$ as their “standard” counterparts and also preserve the linear stability of the system if the following conditions are met:
  • Let us assume that for the “standard” numerical method, the linear stability and positivity are preserved if $\Delta t<\varphi^{*}$. Then, the nonstandard method preserves these properties if $\varphi(\Delta t)<\varphi^{*}$.
  • $\varphi(\Delta t)=\Delta t+O((\Delta t)^{p+1})$.
Note that both of these conditions are met if we choose a function $\varphi$ by the ideas discussed in Section 3.3 with $\mathcal{B}=\varphi^{*}$.
Moreover, in the next Theorems we are always assuming that the one-step methods (that are used to calculate the initial values) preserve a given qualitative property. By using the strongly stability preserving form of “standard” Runge-Kutta methods, conditions under which these requirements are fulfilled can be obtained (see for example [61]). Then, if we replace the usual stepsize $\Delta t$ by $\varphi(\Delta t)$ in these one-step methods, the required properties can be guaranteed for the nonstandard versions of these methods for any stepsize $\Delta t>0$.
@@
We also consider the discrete version of the same property - namely, we say that the numerical solution $u^{n}=(u_{1}^{n},\;u_{2}^{n},\;\dots,\;u_{m}^{n})^{T}$ preserves the boundedness (from above) for element $u_{j}^{n}$, if in the case of $u_{j}^{0}\leq M$ we have $u_{j}^{n}\leq M$ for every $n=1,\;2,\;\dots,\;\mathcal{N}$. Again, a similar definition can also be made for boundedness from below.
@@
Another theorem can be stated if we consider the monotone property of the solution. Let us assume that the component $y_{k}(t)$ of the solution of equation (1) is monotone increasing for $t\in[t_{0},T]$ if $\tilde{y}\in(A,B)$ element-wise. Here $A,B\in\mathbb{R},\;A<B$, where we also allow that $A=$-$\infty$ or $B=\infty$, meaning that we also allow cases like $\tilde{y}\in[0,\infty)$ and not only $\tilde{y}\in[0,C]$ ($C>0$).
@@
  • For the starting values of the multistep method (10), we have
    $u_{k}^{0},\;u_{k}^{1},\;\dots,\;u_{k}^{s-1}\in(A,B).$
@@
Note that property (11) is only a necessary, but not sufficient condition for the monotone increase of numerical solution $u_{k}^{n}$. Numerical experiments indicate that the numerical scheme can indeed preserve the “classical” monotone property (i.e. $u_{k}^{n}\leq u_{k}^{n+1}$) if the bound for $\varphi(x)$ is even smaller than $\mathcal{C}B_{FE}$, but if we only assume that $\varphi(x)\leq\mathcal{C}B_{FE}$, then only inequality (11) holds (see the tests in Section 5). The determination of a bound under which the “classical monotone property” is also preserved could be a topic of further research.
@@
  • For the starting values of the multistep method (10), we have
    $P(u_{1}^{n},\;u_{2}^{n},\;\dots,u_{m}^{n})=\mathcal{M}$
    and $u^{n}\in S$ holds for $n=0,\;1,\;\dots,\;s-1$.